# EnterpriseBench Corecraft: Training Generalizable Agents on High-Fidelity RL Environments

### v2

@@ Abstract
We show that training AI agents on high-fidelity reinforcement learning environments produces capabilities that generalize beyond the training distribution. We introduce Corecraft, the first environment in EnterpriseBench, Surge AI’s suite of agentic RL environments. Corecraft is a fully operational enterprise simulation of a customer support organization, comprising over 2,500 entities across 14 entity types with 23 unique tools, designed to measure whether AI agents can perform the multi-step, domain-specific work that real jobs demand. Frontier models such as GPT-5.2 and Claude Opus 4.6 solve fewer than 30% of tasks when all expert-authored rubric criteria must be satisfied. Using this environment, we train GLM 4.6 with Group Relative Policy Optimization (GRPO) and adaptive clipping. After a single epoch of training, the model improves from 25.37% to 36.76% task pass rate on held-out evaluation tasks. More importantly, these gains transfer to out-of-distribution benchmarks: +4.5% on BFCL Parallel, +7.4% on $\tau^{2}$-Bench Retail, and +6.8% on Tool Decathlon (Pass@1). We believe three environment properties are consistent with the observed transfer: task-centric world building that optimizes for diverse, challenging tasks; expert-authored rubrics enabling reliable reward computation; and enterprise workflows that reflect realistic professional patterns. Our results suggest that environment quality, diversity, and realism are key factors enabling generalizable agent capabilities.
@@ 4 Frontier Model Performance on Corecraft
To characterize the difficulty of Corecraft and establish the headroom available for training, we evaluated frontier models on the current Corecraft benchmark task set. Table 1 reports task pass rates, where a task is counted as passed only when all rubric criteria are satisfied across trajectories.
@@
GPT-5.2 High 29.20
Claude Opus 4.6 (Adaptive) 26.00
DeepSeek-V3.2 Thinking 23.10
Claude Opus 4.6 (Standard) 22.10
GPT-5.2 Codex xHigh 20.10
Grok-4.1 Fast Reasoning 20.00
Gemini 3 Flash 18.00
@@
Gemini 3 Pro 13.00
Nova 2 Pro 8.90
Kimi-K2.5 8.70
Qwen3-Max Thinking 3.60
@@
Llama 4 Maverick 0.00
Table 1: Frontier model performance on the current Corecraft benchmark task set. Pass rate requires all rubric criteria to be satisfied. Even the best model (GPT-5.2 High) solves fewer than 30% of tasks.
These results reveal substantial headroom. Even GPT-5.2, the strongest model on this set, achieves only 29.20% when all rubric criteria must be satisfied. This shows that Corecraft presents challenging tasks that expose capability limitations even in the most advanced models, providing a good training signal for RL.
@@ 9 Acknowledgements
We thank the Surge AI workforce for their essential contributions to this research. Domain experts designed workplace tasks with complexity and realistic edge cases that reflect actual operational challenges. Annotators populated the Corecraft environment with coherent entities, relationships, and data that approximate real-world e-commerce systems. Expert evaluators assessed model trajectories and identified failure patterns, enabling the systematic analysis presented in this work. We thank Ian Niebres for setting up the Corecraft environment, Tim Bauman for leading Corecraft engineering and John Stacy for conducting evaluations and analyses on external benchmarks.

### v3

@@ Abstract
We show that training AI agents on high-fidelity reinforcement learning environments produces capabilities that generalize beyond the training distribution. We introduce Corecraft, the first environment in EnterpriseBench, Surge AI’s suite of agentic RL environments. Corecraft is a fully operational enterprise simulation of a customer support organization, comprising over 2,500 entities across 14 entity types with 23 unique tools, designed to measure whether AI agents can perform the multi-step, domain-specific work that real jobs demand. Frontier models such as Claude Opus 4.6, GPT-5.2, and Gemini 3.1 Pro solve fewer than 35% of tasks when all expert-authored rubric criteria must be satisfied. Using this environment, we train GLM 4.6 with Group Relative Policy Optimization (GRPO) and adaptive clipping. After a single epoch of training, the model improves from 25.37% to 36.76% task pass rate on held-out evaluation tasks. More importantly, these gains transfer to out-of-distribution benchmarks: +4.5% on BFCL Parallel, +7.4% on $\tau^{2}$-Bench Retail, and +6.8% on Tool Decathlon (Pass@1). We believe three environment properties are consistent with the observed transfer: task-centric world building that optimizes for diverse, challenging tasks; expert-authored rubrics enabling reliable reward computation; and enterprise workflows that reflect realistic professional patterns. Our results suggest that environment quality, diversity, and realism are key factors enabling generalizable agent capabilities.
@@ 4 Frontier Model Performance on Corecraft
To characterize the difficulty of Corecraft and establish the headroom available for training, we evaluated frontier models on the current Corecraft benchmark task set. Table 1 reports task pass rates, where a task is counted as passed only when all rubric criteria are satisfied across trajectories. These results reveal substantial headroom and even the strongest model achieves only 30.80% task pass rate on this set.
@@
Claude Opus 4.6 (Adaptive Thinking + Max Reasoning Effort) 30.80
GPT-5.2 (High Reasoning Effort) 29.70
Gemini 3.1 Pro 27.20
Claude Opus 4.6 (High Reasoning Effort) 26.20
DeepSeek v3.2 Thinking 24.10
Claude Opus 4.6 22.10
Grok 4.1 Fast Reasoning 20.50
GPT-5.2-Codex (xHigh Reasoning Effort) 20.10
Gemini 3 Flash 20.00
@@
Claude Sonnet 4.6 16.40
Gemini 3 Pro 14.40
Qwen 3.5 Plus 11.3
Nova 2 Pro (High Reasoning Effort) 8.90
Kimi K2.5 8.70
@@
Qwen3 Max Thinking 3.60
Table 1: Frontier model performance on the current Corecraft benchmark task set. Pass rate requires all rubric criteria to be satisfied. Even the best model (Claude Opus 4.6 Adaptive) solves fewer than 31% of tasks.
@@ 9 Acknowledgements
We thank the Surge AI workforce for their essential contributions to this research. Domain experts designed workplace tasks with complexity and realistic edge cases that reflect actual operational challenges. Annotators populated the Corecraft environment with coherent entities, relationships, and data that approximate real-world e-commerce systems. Expert evaluators assessed model trajectories and identified failure patterns, enabling the systematic analysis presented in this work. We thank Tim Bauman for leading Corecraft engineering and John Stacy for conducting thorough evaluations and analyses on external benchmarks.